OpenStudy (anonymous):

Chapter 5:Indices and Logarithms Solve the following equation @ganeshie8 @ikram002p

OpenStudy (anonymous):

\[4^{n+3}=7^{n-1}\]

ganeshie8 (ganeshie8):

take natural log both sides

ganeshie8 (ganeshie8):

then use below logarithm property : \[\large \ln (a^{\color{red}{x}}) = \color{Red}{x}\times \ln (a)\]

OpenStudy (anonymous):

\[\log _{10}4^{n+3}=\log _{10}7^{n-1}\]

OpenStudy (anonymous):

\[(n+3)~\times~\log _{10}4=(n-1)~\times~\log _{10}7\]

OpenStudy (anonymous):

what should we do next?

ganeshie8 (ganeshie8):

distribute..

ganeshie8 (ganeshie8):

\[(\color{Red}{n}+3)~\times~\log _{10}4=(\color{Red}{n}-1)~\times~\log _{10}7\] \[\color{Red}{n}\log _{10}4 + 3\log _{10}4=\color{Red}{n}\log _{10}7- \log_{10} 7\]

ganeshie8 (ganeshie8):

collect terms with \(\color{Red}{n}\) on one side and factor \(\color{Red}{n}\)

OpenStudy (anonymous):

\[n~\log _{10}4-n~\log _{10}7=-3\log _{10}4-\log _{10}7\]

OpenStudy (anonymous):

\[n(\log _{10}4-\log _{10}7)=-3\log _{10}4-\log _{10}7\]

OpenStudy (anonymous):

\[n =\frac{ -3\log _{10}4-\log _{10}7 }{ \log _{10}4-\log _{10}7 }\]

OpenStudy (anonymous):

n=10.91

ganeshie8 (ganeshie8):

Yes!

OpenStudy (anonymous):

Thnx @ganeshie8